### v3

# Controlling Continuous Relaxation for Combinatorial Optimization

Abstract
Unsupervised learning (UL)-based solvers for combinatorial optimization (CO) train a neural network whose output provides a soft solution by directly optimizing the CO objective using a continuous relaxation strategy. These solvers offer several advantages over traditional methods and other learning-based methods, particularly for large-scale CO problems. However, UL-based solvers face two practical issues: (I) an optimization issue where UL-based solvers are easily trapped at local optima, and (II) a rounding issue where UL-based solvers require artificial post-learning rounding from the continuous space back to the original discrete space, undermining the robustness of the results. This study proposes a Continuous Relaxation Annealing (CRA) strategy, an effective rounding-free learning method for UL-based solvers. CRA introduces a penalty term that dynamically shifts from prioritizing continuous solutions, effectively smoothing the non-convexity of the objective function, to enforcing discreteness, eliminating the artificial rounding. Experimental results demonstrate that CRA significantly enhances the performance of UL-based solvers, outperforming existing UL-based solvers and greedy algorithms in complex CO problems. It also effectively eliminates the artificial rounding and accelerates the learning.
1 Introduction
The objective of combinatorial optimization (CO) problems is to find the optimal solution from a discrete space, and these problems are fundamental in many real-world applications (Papadimitriou and Steiglitz, 1998). Most CO problems are NP-hard or NP-complete, making it challenging to solve large-scale problems within feasible computational time. Traditional methods frequently depend on heuristics to find approximate solutions; however, considerable insights into the specific problems are required. Recently, several studies have used machine learning methods to handle CO problems by learning heuristics. Most of these studies focus on supervised learning (SL)-based solvers (Hudson et al., 2021; Joshi et al., 2019; Gasse et al., 2019; Selsam et al., 2018; Khalil et al., 2016), which require optimal solutions to the CO problem as supervision during training. However, optimal solutions are hard to get in practice, and SL-based solvers often do not generalize well (Yehuda et al., 2020). Reinforcement learning (RL)-based solvers (Yao et al., 2019; Chen and Tian, 2019; Yolcu and Póczos, 2019; Nazari et al., 2018; Khalil et al., 2017; Bello et al., 2016) do not require such optimal solutions but frequently suffer from notoriously unstable training due to poor gradient estimation and hard explorations (Mnih et al., 2015; Tang et al., 2017; Espeholt et al., 2018). Recently, unsupervised learning (UL)-based solvers have attracted much attention (Schuetz et al., 2022a; Karalias and Loukas, 2020; Amizadeh et al., 2018). UL-based solvers follow a continuous relaxation approach. They train a UL model to output a “soft” solution to the relaxed CO problem by directly optimizing the differentiable objective function, offering significantly stable and fast training even for large-scale CO problems. Notably, physics-inspired GNN (PI-GNN) solver (Schuetz et al., 2022a) employs graph neural networks (GNN) to automatically learn specific heuristics for a single instance and performs on par with or outperforms existing solvers for CO problems with millions of variables without optimal solutions.
While offering some advantages over traditional and other machine learning-based solvers, they face two practical issues. The first issue is an optimization issues where UL-based solvers are easily trapped at local optima. Angelini and Ricci-Tersenghi (2023) demonstrated that the PI-GNN solver could not achieve results comparable to those of the degree-based greedy algorithm (DGA) (Angelini and Ricci-Tersenghi, 2019) in maximum independent set (MIS) problems on random regular graphs (RRG). Wang and Li (2023) also pointed out the importance of using dataset/history and initialization of GNN using output of greedy solvers for PI-GNN solver to overcome the optimization issues. This issue is a crucial bottleneck for its applicability in various real-world applications. The second issue relates to the inherent ambiguity in the continuous relaxation approach. This approach requires artificial rounding from the soft solution, which may include continuous values, back to the original discrete solution, potentially undermining the robustness of the results. Indeed, while linear relaxation can provide an optimal solution for original discrete problems on bipartite graphs (Hoffman and Kruskal, 2010), it typically leads to solutions with $\nicefrac{{1}}{{2}}$ values, which is known to half-integrality (Nemhauser and Trotter Jr, 1974), where existing rounding methods (Schuetz et al., 2022b; Wang et al., 2022) completely lose their robustness.
To address these issues, we propose the Continuous Relaxation Annealing (CRA) strategy. The CRA introduces a penalty term to control the continuity and discreteness of the relaxed variables, with a parameter $\gamma$ to regulate the intensity of this penalty term. If the parameter $\gamma$ is small, the relaxed variable tends to favor continuous solutions. In contrast, a large $\gamma$ value leads the relaxed variables to favor the discrete space. This penalty term also effectively eliminates local optimum. Moreover, a small $\gamma$ forces the loss function to approach a simple convex function, encouraging active exploration in the continuous space. The CRA also includes an annealing process, where $\gamma$ is increased progressively until the relaxed variables are nearly discrete values, eliminating the artificial rounding from the continuous to the original discrete space after learning. In this study, the solver that applies the CRA to the PI-GNN solver is referred to as the CRA-PI-GNN solver.
We also demonstrate the benefits of the CRA through experiments on benchmark CO problems, including MIS, maximum cut (MaxCut), and diverse bipartite matching (DBM) problems across graphs of varying sizes and degrees. The experimental results show that the CRA significantly enhances the performance of the PI-GNN solver, outperforming the original PI-GNN solver, other state-of-the-art learning-based baselines, and greedy algorithms. This improvement is achieved by directly optimizing each instance without any history/dataset. Additionally, these experiments indicate that the CRA accelerates the learning process of the PI-GNN solver. Notably, these results overcome the limitations pointed out by Angelini and Ricci-Tersenghi (2023); Wang and Li (2023), highlighting the further potential of UL-based solvers.
Notation
We use the shorthand expression $[N]=\{1,2,\ldots,N\}$,  $N\in{\mathbb{N}}$. Here, $I_{N}\in{\mathbb{R}}^{N\times N}$ denotes an $N\times N$ identity matrix, ${\bm{1}}_{N}$ denotes the vector $(1,\ldots,1)^{\top}\in{\mathbb{R}}^{N}$, and ${\bm{0}}_{N}$ denotes the vector $(0,\ldots,0)^{\top}\in{\mathbb{R}}^{N}$. $G(V,E)$ denotes an undirected graph, where $V$ is the set of nodes with cardinality $|V|=N$, and $E\subseteq V\times V$ is the set of edges. For a graph $G(V,E)$, $A_{ij}$ denote the adjacency matrix with $A_{ij}=0$ if an edge $(i,j)$ does not exist and $A_{ij}>0$ if the edge connects.
2 Background
Combinatorial optimization
The goal of this study is to solve the following CO problem.
$\min_{{\bm{x}}\in\{0,1\}^{N}}f({\bm{x}};C)~{}~{}~{}\mathrm{s.t.}~{}~{}~{}{\bm{x}}\in{\mathcal{X}}(C),$ (1)
where $C\in{\mathcal{C}}$ denotes instance parameters, e.g., a graph $G=(V,E)$, where $V$ is the node set, and $E$ is the edge set, and $f:{\mathcal{X}}\times{\mathcal{C}}\to{\mathbb{R}}$ represents the cost function. In addition, ${\bm{x}}=(x_{i})_{1\leq i\leq N}\in\{0,1\}^{N}$ is a binary vector to be optimized, and ${\mathcal{X}}(C)\subseteq\{0,1\}^{N}$ is the space of feasible solutions, typically specified by the following equality and inequality constraints,
${\mathcal{X}}(C)=\{{\bm{x}}\in\{0,1\}^{N}\mid\forall i\in[I],~{}g_{i}({\bm{x}};C)\leq 0,~{}\forall j\in[J],~{}h_{j}({\bm{x}};C)=0\}.$
Here, for $i\in[I]$, $g_{i}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the inequality constraint, and for $j\in[J]$, $h_{j}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the equality constraint. Following UL-based solvers (Wang et al., 2022; Schuetz et al., 2022a; Karalias and Loukas, 2020), we first rewrite the constrained problem into an equivalent unconstrained form via the penalty method (Smith et al., 1997):
[…]
Here, for all $i\in[I+J]$, $v_{i}:\{0,1\}^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the penalty term, which increases the value when the constraints are violated, e.g., the penalty term is defined as follows:
$\forall i\in[I],j\in[J],~{}v_{i}({\bm{x}};C)=\max(0,g_{i}({\bm{x}};C)),~{}~{}\forall j\in[J],~{}v_{j}({\bm{x}};C)=(h_{j}({\bm{x}};C))^{2},$
and ${\bm{\lambda}}=(\lambda_{i})_{1\leq i\leq I+J}\in{\mathbb{R}}^{I+J}$ represents the penalty parameters that control the tradeoff between satisfying the constraints and optimizing the cost. Note that the penalty for violations of the constraints becomes more significant as ${\bm{\lambda}}$ increases. In the following, we show the example of this formulation.
Example: MIS problem
The MIS problem is a fundamental NP-hard problem (Karp, 2010) defined as follows. Given an undirected graph $G(V,E)$, an independent set (IS) is a subset of nodes ${\mathcal{I}}\in V$ where any two nodes in the set are not adjacent. The MIS problem attempts to find the largest IS, which is denoted ${\mathcal{I}}^{\ast}$. In this study, $\rho$ denotes the IS density, where $\rho=|{\mathcal{I}}|/|V|$. To formulate the problem, a binary variable $x_{i}$ is assigned to each node $i\in V$. Then the MIS problem is formulated as follows:
$f({\bm{x}};G,\lambda)=-\sum_{i\in V}x_{i}+\lambda\sum_{(i,j)\in E}x_{i}x_{j},$ (3)
where the first term attempts to maximize the number of nodes assigned $1$, and the second term penalizes the adjacent nodes marked $1$ according to the penalty parameter $\lambda$.
2.1 Unsupervised learning based solvers
Learning for CO problems involve leaning an algorithm ${\mathcal{A}}_{{\bm{\theta}}}(\cdot):{\mathcal{C}}\to\{0,1\}^{N}$ parameterized by a nural network (NN), where ${\bm{\theta}}$ denotes the parameters. Given an instance parameter $C\in{\mathcal{C}}$, this algorithm generates a valid solution $\hat{{\bm{x}}}={\mathcal{A}}_{\theta}(C)\in{\mathcal{X}}(C)$ and minimizes $f(\hat{{\bm{x}}};C)$. Several approaches have been proposed to learn ${\mathcal{A}}_{\theta}$. This study focuses on UL-based solvers, where no labeled solution ${\bm{x}}^{\ast}\in\mathrm{argmin}_{{\bm{x}}\in{\mathcal{X}}(C)}f({\bm{x}};C)$ is employed during training (Wang et al., 2022; Schuetz et al., 2022a; Karalias and Loukas, 2020; Amizadeh et al., 2018). In the following, we describe the details of the UL-based solvers. The UL-based solvers employ continuous relaxation strategy for directly training neural networks (NN). The continuous relaxation strategy involves reformulating an original CO problem into a continuous optimization problem by converting discrete variables into continuous variables. A typical example of continuous relaxation can be expressed as follows:
[…]
where ${\bm{p}}=(p_{i})_{1\leq i\leq N}\in[0,1]^{N}$ represents a set of relaxed continuous variables, i.e., each binary variable $x_{i}\in\{0,1\}$ is relaxed to a continuous one $p_{i}\in[0,1]$, and $\hat{f}:[0,1]^{N}\times{\mathcal{C}}\to{\mathbb{R}}$ is the relaxation of $f$, satisfying $\hat{f}({\bm{x}};C)=f({\bm{x}};C)$ for ${\bm{x}}\in\{0,1\}^{N}$. The relation between the constraint $v_{i}$ and its relaxation $\hat{v}_{i}$ is similar for $i\in[I+J]$, i.e., $\forall i\in[I+J],~{}\hat{v}_{i}({\bm{x}};C)=v_{i}({\bm{x}};C)$ for ${\bm{x}}\in\{0,1\}^{N}$. Wang et al. (2022) and Schuetz et al. (2022a) formulated ${\mathcal{A}}_{\theta}(C)$ as the relaxed continuous variables, i.e., ${\mathcal{A}}_{\theta}(\cdot):{\mathcal{C}}\to[0,1]^{n}$. In the following discussions, to represent the parametrization of the relaxed variables explicitly, we denote ${\mathcal{A}}_{\theta}$ by ${\bm{p}}_{\theta}$. Then, ${\bm{p}}_{\theta}$ is optimized by minimizing the following label-independent function directly:
[…]
After training the relaxed solution ${\bm{p}}_{{\bm{\theta}}}$ is transformed into discrete variables by artificial rounding ${\bm{p}}_{{\bm{\theta}}}$, such that $\forall i\in[N],~{}x_{i}=\mathrm{int}(p_{{\bm{\theta}},i}(C))$ using a threshold (Schuetz et al., 2022a), or via a greedy method (Wang et al., 2022). Based on this formulation, two types of schemes for UL-based solvers have been developed:
(Type I) Learning generalized heuristics from history/data
One approach, proposed by Karalias and Loukas (2020), aims to automatically learn typically effective heuristics from history/dataset instances ${\mathcal{D}}=\{C_{\mu}\}_{\mu=1}^{P}$ and then apply these learned heuristics to new instance $C^{\ast}$ through inference. Note that this method assumes that the training dataset can be easily obtained or that meaningful data augmentation is possible. Specifically, given a set of training instances ${\mathcal{D}}=(C_{\mu})$, IID sampled from a distribution $P(C)$, the goal is to minimize the averaged loss function $\min_{{\bm{\theta}}}\sum_{\mu=1}^{P}l({\bm{\theta}};C_{\mu},{\bm{\lambda}})$. However, this method cannot guarantee quality for a test instance $C^{\ast}$. Even if the training instances ${\mathcal{D}}$ are numerous and the test instance $C$ also follows $P(C)$, a low averaged performance ${\mathbb{E}}_{C\sim P(C)}[\hat{l}(\theta;C)]$ may not guarantee a low $l(\theta;C)$ for on specific $C$. To address this issues, Wang and Li (2023) introduced a meta-learning approach, where NNs aim for good initialization for future instances rather than direct solutions.
(Type II) Learning effective heuristics on a specific single instance
Another approach, referred to as the PI-GNN solver (Schuetz et al., 2022a, b), automatically learns instance-specific heuristics for a given single instance using the instance parameter $C$ by directly employing Eq. (4). This approach studied CO problems on graphs, i.e., $C=G(V,E)$, and employed GNNs for the relaxed variables $p_{\theta}(G)$. An $L$-layered GNN is trained to directly minimize $\hat{l}({\bm{\theta}};C,{\bm{\lambda}})$ in Eq. (4), taking as input a graph $G$ and the embedding vectors on the nodes, and outputting the relaxed solution ${\bm{p}}_{{\bm{\theta}}}(G)\in[0,1]^{N}$. A detailed description of GNNs is given in Appendix E.2. Note that this setting is applicable even when the training dataset ${\mathcal{D}}$ is not easily obtainable. However, learning to minimize Eq. 4 for a single instance can be time-consuming compared with the inference process. Nonetheless, for large-scale CO problems, Schuetz et al. (2022a, b) demonstrated superiority over other solvers in terms of both time and solution performance. This study focuses on the advancement of this UL based solver for a single instance.
3 Continuous relaxation annealing for UL-based solvers
In this section, we first discuss the practical issues of UL-based solvers and then introduce continuous relaxation annealing (CRA) as a solution to these issues.
3.1 Motivation: practical issues of UL-based solvers
UL-based solvers, for a single instance CO problems, are effective for addressing large-scale CO problems. However, they present following two practical issues, indicated in several previous studies (Wang and Li, 2023; Angelini and Ricci-Tersenghi, 2023). In addition, we also numerically validate these issues; see Appendix F.1 for detailed results.
(i) Ambiguity in rounding method after learning
UL-based solvers employ a continuous relaxation strategy to train NNs and then convert the relaxed continuous variables into discrete binary variables through artificial rounding as discussed in Section 2.1. This inherent ambiguity in the continuous relaxation strategy often leads to potential discrepancies between the optimal solutions of the original discrete CO problem and those of the relaxed continuous versions. The continuous relaxation expands the solution space, often leading to the emergence of continuous values that reduce the cost compared to an optimal binary value. Indeed, while linear relaxation can provide an optimal solution for original discrete problems on bipartite graphs (Hoffman and Kruskal, 2010), it typically leads to solutions with $\nicefrac{{1}}{{2}}$ values, which is known to half-integrality (Nemhauser and Trotter Jr, 1974). Existing rounding methods lose their robustness in these scenarios. In deed, PI-GNN solver often outputs values close to $\nicefrac{{1}}{{2}}$, demonstrating the limitations of current rounding techniques for UL-based solvers.
(ii) Difficulty in optimizing NNs
Recently, Angelini and Ricci-Tersenghi (2023) demonstrated that PI-GNN solver could not achieve results comparable to those of the degree-based greedy algorithm (DGA) (Angelini and Ricci-Tersenghi, 2019) in solving the MIS problems on RRGs. Angelini and Ricci-Tersenghi (2023) also emphasized the importance of evaluating UL-based solvers in complex CO problems, where the performance of greedy algorithms worsens. The representative example is MIS problems on RRGs with a constant degree $d>16$, where a clustering transition occurs in the solution space, forming barriers that impede optimization. Moreover, Wang and Li (2023) emphasized the importance of using training/historical datasets, ${\mathcal{D}}=\{C_{\mu}\}_{1\leq\mu\leq P}$, which contains various graphs and initialization using the outputs of greedy solvers, e.g., DGA and RGA for MIS problems. Their numerical analysis indicated that PI-GNN solver tends to be trapped in local optima when directly optimized for a single instance without leveraging such a training dataset. However, in a practical setting, systematic methods for generating or collecting training datasets, ${\mathcal{D}}$, that effectively avoid local optima remains unclear. The extra training for instances that do not contribute to escaping local optima is also time-consuming. Thus, developing an effective UL-based solver for a single instance is crucial without relying on training data ${\mathcal{D}}$. Our numerical experiments, detailed in Appendix F.1, also validated this optimization issue, demonstrating that as the problem complexity increases, PI-GNN solver tends to be absorbed into the trivial local optima ${\bm{p}}_{{\bm{\theta}}}={\bm{0}}_{N}$ in some problems, resulting in prolonged plateaus which significantly decelerates the learning process and, in particularly challenging cases, can render learning entirely infeasible.
3.2 Continuous relaxation annealing
Penalty term to control discreteness and continuity
To address these issues, we propose a penalty term to control the discreteness and continuity of the relaxed variables, formulated as follows:
[…]
Theorem 3.1.
Under the assumption that the objective function $\hat{l}({\bm{p}};C)$ is bounded within the domain $[0,1]^{N}$, as $\gamma\to+\infty$, the soft solutions ${\bm{p}}^{\ast}\in\mathrm{argmin}_{{\bm{p}}}\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$ converge to the original solutions ${\bm{x}}^{\ast}\in\mathrm{argmin}_{{\bm{x}}}l({\bm{x}};C,{\bm{\lambda}})$. In addition, as $\gamma\to-\infty$, the loss function $\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$ becomes convex, and the soft solution $\nicefrac{{{\bm{1}}_{N}}}{{2}}=\mathrm{argmin}_{{\bm{p}}}\hat{r}({\bm{p}},C,{\bm{\lambda}},\gamma)$ is unique.
Refer to Appendix B.1 for the detailed proof. Theorem 3.1 can be generalized for any convex function $\Phi({\bm{p}};C)$ which has a unique maximum at $\nicefrac{{{\bm{1}}_{N}}}{{2}}$ and becomes a global minimum for all ${\bm{p}}\in\{0,1\}^{N}$, e.g., binary cross entropy $\Phi_{\mathrm{CE}}({\bm{p}})=\sum_{i=1}^{N}(p_{i}\log p_{i}+(1-p_{i})\log(1-p_{i}))$ introduced in Sun et al. (2022); Sanokowski et al. (2024) for the UL-based solvers using history/data. Additionally, the penalty term eliminates the stationary point ${\bm{p}}^{\ast}={\bm{0}}_{n}$ in Section 3.1, avoiding convergence to the plateau. For UL-based solvers, the penalty term is expressed as follows:
$\hat{r}({\bm{\theta}};C,{\bm{\lambda}},\gamma)=\hat{l}({\bm{\theta}};C,{\bm{\lambda}})+\gamma\Phi({\bm{\theta}};C),$ (6)
[…]
Annealing penalty term
[…]
Figure 1: The value $\gamma<0$ facilitates exploration by smoothing the objective function. As $\gamma$ increases, it enforces optimal discrete solutions by smoothing out suboptimal continuous ones.
On the other hand, the gradient of our penalty term in Eq. 5, $\nicefrac{{\partial\Phi({\bm{p}})}}{{\partial p_{i}}}$, is bounded within the range of $[-2\alpha,2\alpha]$ for any $\gamma$, preventing the gradient divergence issue in $\Phi_{\mathrm{CE}}({\bm{p}})$. Additionally, by increasing $\alpha$, the absolute value of the gradient near $\nicefrac{{1}}{{2}}$ becomes smaller, allowing control of the strength of smoothing out towards the discrete solution near $\nicefrac{{1}}{{2}}$. We also propose an early stopping strategy that monitors the loss function and the penalty term. The annealing and learning processes are stopped when the penalty term approaches zero, i.e., $\Phi({\bm{\theta}};C)\approx 0$. Various annealing schedules can be considered; This study employ the following scheduling strategy: $\gamma(\tau+1)\leftarrow\gamma(\tau)+\varepsilon$, where the scheduling rate $\varepsilon\in{\mathbb{R}}_{+}$ is a small constant, and $\tau$ denotes the update iterations of the trainable parameters. We refer to the PI-GNN solver with this continuous relaxation annealing as CRA-PI-GNN solver. Here, two additional hyperparameters are introduced: the initial scheduling value $\gamma(0)$ and the scheduling rate $\varepsilon$. Our numerical experiments suggest that better solutions are obtained when $\gamma(0)$ is set to a small negative value and $\varepsilon$ is kept small; the ablation study for these parameters in Appendix F.3.
4 Related Work
Previous works on UL-based solvers have addressed various problems, such as MaxCut problems (Yao et al., 2019) and traveling salesman problems (Hudson et al., 2021) relying on carefully designed problem-specific objectives. Some studies have also investigated constraint satisfaction problems (Amizadeh et al., 2018; Toenshoff et al., 2019), but applying these approaches to general CO problems typically requires problem reductions. Karalias and Loukas (2020) proposed Erdős Goes Neural (EGN) solver, unsupervised learning framework for general CO problems based on Erdős’ probabilistic method. This solver can generate solutions through inference process using training instances. Subsequently, Wang et al. (2022) proposed an entry-wise concave continuous relaxation, which generalized EGN solver to wider CO problems. On the other hand, Schuetz et al. (2022a, b) proposed PI-GNN solver, UL-based solver for a single instance, automatically learning problem-specific heuristics through training process. However, Angelini and Ricci-Tersenghi (2023); Boettcher (2023) pointed out the optimization difficulties where PI-GNN solver could not achieve results comparable to those of greedy algorithms. Wang and Li (2023) also claimed optimization issues, even for the EGN solver, and emphasized the importance of learning from training data/history to overcome local optima and then proposed Meta-EGN solvers, a meta-learning approach that updates NN network parameters for individual CO problem instances. Then, Wang and Li (2023) introduced a meta-learning method for EGN solvers, which NN parameters are updated for individual CO problem instances. To address these optimization issue, Sun et al. (2022); Sanokowski et al. (2024) proposed annealing strategy similar to simulated annealing (Kirkpatrick et al., 1983).
5 Experiments
We begin by evaluating the performance of CRA-PI-GNN solver on the MIS and the MaxCut benchmark problems across multiple graphs of varying sizes, demonstrating that CRA effectively overcomes optimization challenges without relying on data/history ${\mathcal{D}}=\{C_{\mu}\}$. We then extend the evaluation to the DBM problems, showing the applicability to more practical CO problems. For the objective functions and the detailed explanations, refer to Appendix E.1.
Baseline methods
In all experiments, the baseline methods include PI-GNN solver (Schuetz et al., 2022a) as the direct baseline of a UL-based solver for a single instance. For the MIS problems, we also consider the random greedy algorithm (RGA) and DGA (Angelini and Ricci-Tersenghi, 2019) as heuristic baselines. For the MaxCut problems, RUN-CSP solver (Toenshoff et al., 2019) is considered as an additional baseline, and a standard greedy algorithm and SDP based approximation algorithm (Goemans and Williamson, 1995) are considered as an additional classical baseline. Both are implemented by (Mehta, 2019). Note that we do not consider UL-based solvers for learning generalized heuristics (Karalias and Loukas, 2020; Wang et al., 2022; Wang and Li, 2023), which rely on training instances ${\mathcal{D}}=\{C_{\mu}\}_{\mu=1}^{P}$. The objective of this study is to evaluate whether CRA-PI-GNN solver can surpass the performance of both PI-GNN solver and greedy algorithms without any training dataset. However, for the MIS problem, EGN solver (Karalias and Loukas, 2020) and Meta-EGN solver (Wang and Li, 2023) are considered to confirm that CRA can overcome the optimization issues without training instances.
Implementation
The objective of the numerical experiments is to compare CRA-PI-GNN solver with PI-GNN solver. Thus, we follow the same experimental configuration described as the experiments in Schuetz et al. (2022a), employing a simple two-layer GCV and GraphSAGE (Hamilton et al., 2017) implemented by the Deep Graph Library (DGL) (Wang et al., 2019); Refer to Appendix D.1 for the detailed architectures. We use the AdamW (Kingma and Ba, 2014) optimizer with a learning rate of $\eta=10^{-4}$ and weight decay of $10^{-2}$. The GNNs are trained for up to $5\times 10^{4}$ epochs with early stopping, which monitors the loss function $\hat{l}({\bm{\theta}};{\bm{\lambda}})$ and the penalty term $\Phi({\bm{\theta}};\gamma)$ with tolerance $10^{-5}$ and patience $10^{3}$ epochs; Further details are provided in Appendix D.2. We set the initial scheduling value to $\gamma(0)=-20$ for the MIS and matching problems, and we set $\gamma(0)=-6$ for the MaxCut problems with the scheduling rate $\varepsilon=10^{-3}$ and curve rate $\alpha=2$ in Eq. (6). These values are not necessarily optimal, and refining these parameters can lead to better solutions; Refer to Appendix F.3 and Appendix F.4 for an ablation study of these parameters.
Evaluation metrics
Following Karalias and Loukas (2020), we use the approximation ratio (ApR), formulated as $\mathrm{ApR}=f({\bm{x}};C)/f({\bm{x}}^{\ast};C)$, where ${\bm{x}}^{\ast}$ is optimal solution. For the MIS problems, we evaluate the ApRs using the theoretical optimal cost (Barbier et al., 2013) and the independent set density $\rho$ relative to the theoretical results. For the MaxCut problems on RRGs, we adopt the cut ratio $\nu$ against the theoretical upper bound (Parisi, 1980; Dembo et al., 2017); see Appendix E.1 for the details. All the results for the MIS and MaxCut problems are summarized based on 5 RRGs with different random seeds. In the case of the MaxCut Gset problem, the ApR is calculated compared to the known best cost functions. Regarding the DBM problems, we calculate the ApR against the global optimal, identified using Gurobi 10.0.1 solver with default settings.
5.1 MIS problems
Degree dependency of solutions using CRA
Figure 2: Independent set density of the MIS problem on $d$-RRG. Results for graphs with $N=10{,}000$ nodes (Left) and $N=20{,}000$ nodes (Right). the dashed lines represent the theoretical results (Barbier et al., 2013).
Figure 3: Cut ratio of the MaxCut problem on $d$-RRG as a function of the degree $d$ Results for $N=10{,}000$ (Left) and $N=20{,}000$ (Right). The dashed lines represents the theoretical upper bounds (Parisi, 1980).
First, we compare the performance of the PI-GNN and CRA-PI-GNN solvers using GCV, as in Schuetz et al. (2022a). Fig. 3 shows the independent set density $\rho_{d}$ as a function of degree $d$ obtained by the PI-GNN and CRA-PI-GNN solvers compared with the theoretical results (Barbier et al., 2013). Across all degrees $d$, the CRA-PI-GNN solver outperformed the PI-GNN solver and approached the theoretical results, whereas the PI-GNN solver fail to find valid solutions, especially for $d\geq 15$, as pointed by the previous studies (Angelini and Ricci-Tersenghi, 2023; Wang and Li, 2023).
Response to Angelini and Ricci-Tersenghi (2023) and Wang and Li (2023)
Table 1: ApR in MIS problems on RRGs with $10{,}000$ nodes and node degree $d=20,100$. “CRA” represents the CRA-PI-GNN solver.
Instance Method ApR
20-RRG RGA $0.776\pm 0.001$
DGA $0.891\pm 0.001$
EGN $0.775$ (2023)
META-EGN $0.887$ (2023)
PI-GNN (GCV) $000.000\pm 0$ 0.000
PI-GNN (SAGE) $0.745\pm 0.003$
CRA (GCV) $0.937\pm 0.002$
CRA (SAGE) $\underline{\mathbf{0.963\pm 0.001}}$
100-RRG RGA $0.663\pm 0.001$
DGA $0.848\pm 0.002$
PI-GNN (GCV) $000.000\pm 0$ 0.000
PI-GNN (SAGE) $000.000\pm 0$ 0.000
CRA (GCV) $0.855\pm 0.004$
CRA (SAGE) $\underline{\mathbf{0.924\pm 0.001}}$
Figure 4: (Right) computational runtime (in seconds) of the CRA-PI-GNN solvers with the GraphSage and Conv architectures on $100$-RRG with varying numbers of nodes $N$.Error bars represent the standard deviations of the results.
MIS problems on RRGs with a degree $d$ larger than $16$ is known to be hard problems (Barbier et al., 2013). As discussed in Section 3.1, Angelini and Ricci-Tersenghi (2023); Wang and Li (2023) have posted the optimization concerns on UL-based solvers. However, we call these claim into question by substantially outperforming heuristics DGA and RGA for the MIS on graphs with $d=20,100$, without training/historical instances ${\mathcal{D}}=\{G^{\mu}\}_{\mu=1}^{n}$, as shown in Table 1.
Acceleration of learning speed
Figure 5: The dynamics of cost function for MIS problems on RRGs with $N=10{,}000$ nodes varying degrees $d$ as a function of the number of parameters updates $N_{\mathrm{EPOCH}}$.
We also compared the learning curve between PI-GNN and CRA-PI-GNN solver to confirmed that the CRA-PI-GNN solver does not become trapped in the plateau, ${\bm{p}}_{N}={\bm{0}}_{N}$, as discussed in Section 3.1. Fig. 5 shows the dynamics of the cost functions for the MIS problems with $N=10{,}000$ across $d=3,5,20,100$. Across all degrees, CRA-PI-GNN solver achieves a better solution with fewer epochs than PI-GNN solver. Specifically, PI-GNN solver becomes significantly slower due to getting trapped in the plateau even for graphs with low degrees, such as $d=3,5$. In contrast, CRA-PI-GNN solver can effectively escape from plateaus through the smoothing and automatic rounding of the penalty term when the negative parameter $\gamma>0$.
Computational scaling
We next evaluate the computational scaling of the CRA-PI-GNN solver for MIS problems with large-scale RRGs with a node degree of $100$ in Fig. 1, following previous studies (Schuetz et al., 2022a; Wang and Li, 2023). Fig. 1 demonstrated a moderate super-linear scaling of the total computational time, approximately $\sim N^{1.4}$ for GCN and $\sim N^{1.7}$ for GraphSage. This performance is nearly identical to that of the PI-GNN solver (Schuetz et al., 2022a) for problems on RRGs with lower degrees. It is important note that the runtimes of CRA-PI-GNN solver heavily depend on the optimizer for GNNs and annealing rate $\varepsilon$; thus this scaling remains largely unchanged for problems other than the MIS on $100$ RRG.
5.2 MaxCut problem
Degree dependency of solutions
We first compare the performances of PI-GNN and CRA-PI-GNN solvers with GCV, following Schuetz et al. (2022a). Fig. 3 shows the cut ratio $\nu_{d}$ as a function of the degree $d$ compared to the theoretical upper bound (Parisi, 1980; Dembo et al., 2017). Across all degrees $d$, CRA-PI-GNN solver also outperforms PI-GNN solver, approaching the theoretical upper bound. In contrast, PI-GNN solver fails to find valid solutions for $d>20$ as with the case of the MIS problems in Section 5.1.
Standard MaxCut benchmark test
Following Schuetz et al. (2022a), we next conducted additional experiments on standard MaxCut benchmark instances based on the publicly available Gset dataset (Ye, 2003), which is commonly used to evaluate MaxCut algorithms. Here, we provide benchmark results for seven distinct graphs with thousands of nodes, including Erdös-Renyi graphs with uniform edge probability, graphs in which the connectivity decays gradually from node $1$ to $N$, $4$-regular toroidal graphs, and a very large Gset instance with $N=10{,}000$ nodes. Table 2 shows, across all problems, CRA-PI-GNN solver outperforms both the PI-GNN, RUN-CSP solvers and other greedy algorithm. See Appendix 3 for the results of solving all other Gsets, where consistently, CRA-PI-GNN provides better results as well.
Table 2: ApR for MaxCut on Gset
GRAPH (NODES, EDGES) GREEDY SDP RUN-CSP PI-GNN CRA
G14 ($800$, $4{,}694$) $0.946$ $0.970$ $0.960$ $0.988$ $\underline{\mathbf{0.994}}$
G15 ($800$, $4{,}661$) $0.939$ $0.958$ $0.960$ $0.980$ $\underline{\mathbf{0.992}}$
G22 ($2{,}000$, $19{,}990$) $0.923$ $0.77$ $0.975$ $0.987$ $\underline{\mathbf{0.998}}$
G49 ($3{,}000$, $6{,}000$) $\underline{\mathbf{1.000}}$ $\underline{\mathbf{1,00}}$ $\underline{\mathbf{1.000}}$ $0.986$ $\mathbf{\underline{1.000}}$
G50 ($3{,}000$, $6{,}000$) $\mathbf{\underline{1.000}}$ $\mathbf{\underline{1.000}}$ $\mathbf{\underline{1.000}}$ $0.990$ $\mathbf{\underline{1.000}}$
G55 ($5{,}000$, $12{,}468$) $0.892$- $0.982$ $0.983$ $\underline{\mathbf{0.991}}$
G70 ($10{,}000$, $99{,}9$ 999) $0.886$- $0.970$ $0.982$ $\underline{\mathbf{0.992}}$
Figure 6: ApR on DBM problems.
5.3 Diverse bipartite matching
To evaluate the applicability of the CRA-PI-GNN solver to more practical problems not on graphs, we conducted experiments on DBM problems (Ferber et al., 2020; Mulamba et al., 2020; Mandi et al., 2022); refer to Appendix E.1 for details.
This problems consists of $27$ distinct instances with varying properties, and each instance comprises $100$ nodes representing scientific publications, divided into two groups of $50$ nodes $N_{1}$ and $N_{2}$. The optimization is formulated as follows:
$l({\bm{x}};C,M,{\bm{\lambda}})=-\textstyle\sum_{ij}C_{ij}x_{ij}+\lambda_{1}\textstyle\sum_{i}\mathrm{ReLU}\Big{(}\textstyle\sum_{j}x_{ij}-1\Big{)}+\lambda_{2}\textstyle\sum_{j}\mathrm{ReLU}\Big{(}\sum_{i}x_{ij}-1\Big{)}\\ +\lambda_{3}\mathrm{ReLU}\Big{(}p\textstyle\sum_{ij}x_{ij}-\textstyle\sum_{ij}M_{ij}x_{ij}\Big{)}+\lambda_{4}\mathrm{ReLU}\Big{(}q\sum_{ij}x_{ij}-\sum_{ij}(1-M_{ij})x_{ij}\Big{)},$
where $C\in{\mathbb{R}}^{N_{1}\times N_{2}}$ represents the likelihood of a link between each pair of nodes, an indicator $M_{ij}$ is set to 00 if article $i$ and $j$ share the same subject field ($1$ otherwise) $\forall i\in N_{1}$, and $j\in N_{2}$. The parameters $p,q\in[0,1]$ represent the probability of pairs sharing their field and of unrelated pairs, respectively. As in Mandi et al. (2022), we explore two variations of this problem, with $p=q=$ being $25$% and $5$%, respectively, and these variations are referred to as Matching-1 and Matching-2, respectively. In this experiment, we set $\lambda_{1}=\lambda_{2}=10$ and $\lambda_{3}=\lambda_{4}=25$. Fig 6 shows that the CRA-PI-GNN solver can find better solutions across all instances.
6 Conclusion
This study proposes CRA strategy to address the both optimization issue and rounding issue in UL-based solvers. The CRA strategy introduces a penalty term that dynamically shifts from prioritizing continuous solutions where the non-convexity of the objective function are effectively smoothed, to enforcing discreteness, eliminating the artificial rounding. Experimental results demonstrate that CRA-PI-GNN solver significantly outperforms PI-GNN solver and greedy algorithms across various complex CO problems, including MIS, MaxCut, and DBM problems. CRA approach not only enhances quality of solutions but also accelerates the learning process.
Appendix A Overview
This supplementary material provides extended explanations, implementation details, and additional results.
Appendix B Derivation
B.1 Proof of Theorem 3.1
First, we present three lemmas, and then we demonstrate Theorem 3.1 based on these lemmas.
Lemma B.1.
For any even natural number $\alpha\in\{2n\mid n\in{\mathbb{N}}_{+}\}$, the function $\phi(p)=1-(2p-1)^{\alpha}$ defined on $[0,1]$ achieves its maximum value of $1$ when $p=1/2$ and its minimum value of 00 when $p=0$ or $p=1$.
Proof.
The derivative of $\phi(p)$ relative to $p$ is $\nicefrac{{d\phi(p)}}{{dp}}=-2\alpha(2p-1)$, which is zero when $p=1/2$. This is a point where the function is maximized because the second derivative $\nicefrac{{d^{2}\phi(p)}}{{dp^{2}}}=-4\alpha\leq 0$. In addition, this function is concave and symmetric relative to $p=1/2$ because $\alpha$ is an even natural number, i.e., $\phi(p)=\phi(1-p)$, thereby achieving its minimum value of 00 when $p=0$ or $p=1$. ∎
Lemma B.2.
For any even natural number $\alpha\in\{2n\mid n\in{\mathbb{N}}_{+}\}$, if $\gamma\to+\infty$, minimizing the penalty term $\gamma\Phi({\bm{p}})=\gamma\sum_{i=1}^{N}(1-(2p_{i}-1)^{\alpha})=\gamma\sum_{i=1}^{N}\phi(p_{i})$ enforces that the for all $i\in[N]$, $p_{i}$ is either 00 or $1$ and, if $\gamma\to-\infty$, the penalty term enforces ${\bm{p}}={\bm{1}}_{N}/2$.
Proof.
From Lemma B.1, as $\gamma\to+\infty$, $\phi(p)$ is minimal value when, for any $i\in[N]$, $p_{i}=0$ or $p_{i}=1$. As $\gamma\to-\infty$, $\phi(p;\alpha,\gamma)$ is minimal value when, for any $i\in[N]$, $p_{i}=1/2$. ∎
Lemma B.3.
For any even number $\alpha\in\{2n\mid n\in{\mathbb{N}}_{+}\}$, $\gamma\Phi({\bm{p}})$ is concave when $\lambda>0$ and is a convex function when $\lambda<0$.
Proof.
Note that $\gamma\Phi({\bm{p}})=\gamma\sum_{i=1}^{N}\phi(p_{i})=\gamma\sum_{i=1}^{N}(1-(2p_{i}-1)^{\alpha})$ is separable across its components $p_{i}$. Thus, it is sufficient to prove that each $\gamma\phi_{i}(p_{i};\alpha)$ is concave or convex in $p_{i}$ because the sum of the concave or convex functions is also concave (and vice versa). Thus, we consider the second derivative of $\gamma\phi_{i}(p_{i})$ with respect to $p_{i}$:
$\gamma\frac{d^{2}\phi_{i}(p_{i})}{dp_{i}^{2}}=-4\gamma\alpha.$
If $\gamma>0$, the second derivative is negative for all $p_{i}\in[0,1]$, and this completes the proof that $\gamma\Phi({\bm{p}})$ is a concave function when $\gamma$ is positive (and vice versa). ∎
Combining Lemma B.1, Lemma B.2 and Lemma B.3, one can show the following theorem.
Theorem B.4.
Under the assumption that the objective function $\hat{l}({\bm{p}};C)$ is bounded within the domain $[0,1]^{N}$, as $\gamma\to+\infty$, the soft solutions ${\bm{p}}^{\ast}\in\mathrm{argmin}_{{\bm{p}}}\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$ converge to the original solutions ${\bm{x}}^{\ast}\in\mathrm{argmin}_{{\bm{x}}}l({\bm{x}};C,{\bm{\lambda}})$. In addition, as $\gamma\to-\infty$, the loss function $\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$ becomes convex, and the soft solution $\nicefrac{{{\bm{1}}_{N}}}{{2}}=\mathrm{argmin}_{{\bm{p}}}\hat{r}({\bm{p}},C,{\bm{\lambda}},\gamma)$ is unique.
Proof.
As $\lambda\to+\infty$, the penalty term $\Phi({\bm{p}})$ dominates the loss function $\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$. According to Lemma B.2, this penalty term forces the optimal solution ${\bm{p}}^{\ast}$ to a binary vector whose components, for all $i\in[N]$ $p_{i}^{\ast}$ that are either 00 or $1$ because any non-binary value results in an infinitely large penalty. This effectively restricts the feasible region to the vertices of the unit hypercube, which correspond to the binary vector in $\{0,1\}^{N}$. Thus, as $\lambda\to\infty$, the solutions to the relaxed problem converge to those of the original problem. As $\lambda\to-\infty$, the penalty term $\Phi({\bm{p}})$ also dominates the loss function $\hat{r}({\bm{p}};C,{\bm{\lambda}},\gamma)$ and the $\hat{r}({\bm{p}};C,{\bm{\lambda}})$ convex function from Lemma B.3. According to Lemma B.2, this penalty term forces the optimal solution ${\bm{p}}^{\ast}={\bm{1}}_{N}/2$. ∎
The theorem holds for the cross entropy penalty given by
$\Phi({\bm{p}})=\sum_{i=1}^{N}\left(p_{i}\log(p_{i})+(1-p_{i})\log(1-p_{i})\right)$ (7)
in the UL-based solver using data or history [Sun et al., 2022, Sanokowski et al., 2024] because $\Phi({\bm{p}})$ can similarly satisfy Lemma B.1, Lemma B.2 and Lemma B.3.
Corollary B.5.
Theorem B.4 holds for the following penalty term:
$\Phi({\bm{p}})=\sum_{i=1}^{N}\left(p_{i}\log(p_{i})+(1-p_{i})\log(1-p_{i})\right).$ (8)
Appendix C Generalization of CRA
C.1 Generalization for to Potts variable optimization
This section generalize the penalty term $\Phi({\bm{\theta}};C)$ introduced for binary variables to $K$-Potts variables. $K$-Potts variable is the Kronecker delta $\delta(x_{i},x_{j})$ which equas one whenever $x_{i}=x_{j}$ and zero otherwise and a decision variable takes on $K$ different values, $\forall i\in[N],~{}x_{i}=1,\ldots,K$. For example, graph $K$-coloring problems can be expressed as
$f({\bm{x}};G(V,E))=-\sum_{(i,j)\in E}\delta(x_{i},x_{j})$ (9)
For Potts variables, the output of the GNN is ${\bm{h}}_{{\bm{\theta}}}^{L}=P({\bm{\theta}};C)\in\mathbb{R}^{N\times K}$ and the penalty term can be generalized as follows.
$\Phi({\bm{\theta}};C)=\sum_{i=1}^{N}\left(1-\sum_{k=1}^{K}\left(KP_{i,k}({\bm{\theta}};C)-1\right)^{\alpha}\right).$ (10)
Appendix D Additional implementation details
D.1 Architecture of GNNs
We describe the details of the GNN architectures used in all numerical experiments. The first convolutional layer takes $H_{0}$-dimensional node embedding vectors, ${\bm{h}}_{{\bm{\theta}}}^{0}$ for each node, as input, yielding $H_{1}$-dimensional feature vectors ${\bm{h}}_{{\bm{\theta}}}^{1}$. Then, the ReLU function is applied as a component-wise nonlinear transformation. The second convolutional layer takes the $H_{1}$-dimensional feature vectors, ${\bm{h}}^{1}_{{\bm{\theta}}}$, as input, producing a $H^{(2)}$-dimensional vector ${\bm{h}}^{2}_{{\bm{\theta}}}$. Finally, a sigmoid function is applied to the $H^{(2)}$-dimensional vector ${\bm{h}}^{2}_{{\bm{\theta}}}$, and the output is the soft solution ${\bm{p}}_{{\bm{\theta}}}\in[0,1]^{N}$. As in [Schuetz et al., 2022a], we set $H_{0}=\mathrm{int}(N^{0.8})$ or , $H_{1}=\mathrm{int}(N^{0.8}/2)$ and $H^{2}=1$ for both GCN and GraphSAGE .
D.2 Training settings
For all numerical experiments, we use the AdamW [Kingma and Ba, 2014] optimizer with a learning rate of $\eta=10^{-4}$ and weight decay of $10^{-2}$, and we train the GNNs for up to $10^{5}$ epochs with early stopping set to the absolute tolerance $10^{-5}$ and patience $10^{3}$. As discussed in Schuetz et al. [2022a], the GNNs are initialized with five different random seeds for a single instance because the results are dependent on the initial values of the trainable parameters; thus selecting the best solution.
Appendix E Experiment details
E.1 Benchmark problems
MIS problems
There are some theoretical results for MIS problems on RRGs with the node degree set to $d$, where each node is connected to exactly $d$ other nodes. The MIS problem is a fundamental NP-hard problem [Karp, 2010] defined as follows. Given an undirected graph $G(V,E)$, an independent set (IS) is a subset of nodes ${\mathcal{I}}\in V$ where any two nodes in the set are not adjacent. The MIS problem attempts to find the largest IS, which is denoted ${\mathcal{I}}^{\ast}$. In this study, $\rho$ denotes the IS density, where $\rho=|{\mathcal{I}}|/|V|$. To formulate the problem, a binary variable $x_{i}$ is assigned to each node $i\in V$. Then the MIS problem is formulated as follows:
$f({\bm{x}};G,\lambda)=-\sum_{i\in V}x_{i}+\lambda\sum_{(i,j)\in E}x_{i}x_{j},$ (11)
where the first term attempts to maximize the number of nodes assigned $1$, and the second term penalizes the adjacent nodes marked $1$ according to the penalty parameter $\lambda$. In our numerical experiments, we set $\lambda=2$, following Schuetz et al. [2022a], no violation is observed as in [Schuetz et al., 2022a]. First, for every $d$, a specific value $\rho_{d}^{\ast}$, which is dependent on only the degree $d$, exists such that the independent set density $|{\mathcal{I}}^{\ast}|/|V|$ converges to $\rho_{d}^{\ast}$ with a high probability as $N$ approaches infinity [Bayati et al., 2010]. Second, a statistical mechanical analysis provides the typical MIS density $\rho^{\mathrm{Theory}}_{d}$, as shown in Fig. 3, and we clarify that for $d>16$, the solution space of ${\mathcal{I}}$ undergoes a clustering transition, which is associated with hardness in sampling [Barbier et al., 2013] because the clustering is likely to create relevant barriers that affect any algorithm searching for the MIS ${\mathcal{I}}^{\ast}$. Finally, the hardness is supported by analytical results in a large $d$ limit, which indicates that, while the maximum independent set density is known to have density $\rho^{\ast}_{d\to\infty}=2\log(d)/d$, to the best of our knowledge, there is no known algorithm that can find an independent set density exceeding $\rho^{\mathrm{alg}}_{d\to\infty}=\log(d)/d$ [Coja-Oghlan and Efthymiou, 2015].
MaxCut problems
The MaxCut problem is also a fundamental NP-hard problem [Karp, 2010] with practical application in machine scheduling [Alidaee et al., 1994], image recognition [Neven et al., 2008] and electronic circuit layout design [Deza and Laurent, 1994]. The MaxCut problem is also a fundamental NP-hard problem [Karp, 2010] Given an graph $G=(V,E)$, a cut set ${\mathcal{C}}\in E$ is defined as a subset of the edge set between the node sets dividing $(V_{1},V_{2}\mid V_{1}\cup V_{2}=V,~{}V_{1}\cap V_{2}=\emptyset)$. The MaxCut problems aim to find the maximum cut set, denoted ${\mathcal{C}}^{\ast}$. Here, the cut ratio is defined as $\nu=|\mathcal{C}|/|\mathcal{V}|$, where $|\mathcal{C}|$ is the cardinality of the cut set. To formulate this problem, each node is assigned a binary variable, where $x_{i}=1$ indicates that node $i$ belongs to $V_{1}$, and $x_{i}=0$ indicates that the node belongs to $V_{2}$. Here, $x_{i}+x_{j}-2x_{i}x_{j}=1$ holds if the edge $(i,j)\in{\mathcal{C}}$. As a result, we obtain the following:
$f({\bm{x}};G)=\sum_{i<j}A_{ij}(2x_{i}x_{j}-x_{i}-x_{j}).$ (12)
This study has also focused on the MaxCut problems on $d$-RRGs, for which several theoretical results have been established. Specifically, for each $d$, the maximum cut ratio is given by $\nu_{d}^{\ast}\approx d/4+P_{\ast}\sqrt{d/4}+{\mathcal{O}}(\sqrt{d})$, where $P_{\ast}=0.7632\ldots$ with a high probability as $N$ approaches infinity [Parisi, 1980, Dembo et al., 2017]. Thus, we take $\nu_{d}^{\mathrm{UB}}=d/4+P_{\ast}\sqrt{d/4}$ as an upper bound for the maximum cut ratio in the large $n$ limit.
DBM problems
Here, the topologies are taken from the Cora citation network [Sen et al., 2008], where each node has $1{,}433$ bag-of-words features, and each edge represents likelihood, as predicted by a machine learning model. Mandi et al. [2022] focused on disjoint topologies within the given topology, and they created $27$ distinct instances with varying properties. Each instance comprises $100$ nodes representing scientific publications, divided into two groups of $50$ nodes $N_{1}$ and $N_{2}$. The optimization task is to find the maximum matching, where diversity constraints ensure connections among papers in the same field and between papers of different fields. This is formulated using a penalty method as follows.
$l({\bm{x}};C,M,{\bm{\lambda}})=-\textstyle\sum_{ij}C_{ij}x_{ij}+\lambda_{1}\textstyle\sum_{i}\mathrm{ReLU}\Big{(}\textstyle\sum_{j}x_{ij}-1\Big{)}+\lambda_{2}\textstyle\sum_{j}\mathrm{ReLU}\Big{(}\sum_{i}x_{ij}-1\Big{)}\\ +\lambda_{3}\mathrm{ReLU}\Big{(}p\textstyle\sum_{ij}x_{ij}-\textstyle\sum_{ij}M_{ij}x_{ij}\Big{)}+\lambda_{4}\mathrm{ReLU}\Big{(}q\sum_{ij}x_{ij}-\sum_{ij}(1-M_{ij})x_{ij}\Big{)},$ (13)
where $C\in{\mathbb{R}}^{N_{1}\times N_{2}}$ represents the likelihood of a link between each pair of nodes, an indicator $M_{ij}$ is set to 00 if article $i$ and $j$ share the same subject field ($1$ otherwise) $\forall i\in N_{1}$, and $j\in N_{2}$. The parameters $p,q\in[0,1]$ represent the probability of pairs sharing their field and of unrelated pairs, respectively. As in Mandi et al. [2022], we explore two variations of this problem, with $p=q=$ being $25$% and $5$%, respectively, and these variations are referred to as Matching-1 and Matching-2, respectively. In this experiment, we set $\lambda_{1}=\lambda_{2}=10$ and $\lambda_{3}=\lambda_{4}=25$.
E.2 GNNs
A GNN [Gilmer et al., 2017, Scarselli et al., 2008] is a specialized neural network for representation learning of graph-structured data. GNNs learn a vectorial representation of each node in two steps, i.e., the aggregate and combine steps. The aggregate step employs a permutation-invariant function to generate an aggregated node feature, and in the combine step, the aggregated node feature is passed through a trainable layer to generate a node embedding, known as "message passing" or the "readout phase." Formally, for a given graph $G=(V,E)$, where each node feature ${\bm{h}}^{0}_{v}\in{\mathbb{R}}^{H_{0}}$ is attached to each node $v\in V$, the GNN updates the following two steps iteratively. First, the aggregate step at each $l$-th layer is defined as follows:
${\bm{a}}_{v}^{l}=\mathrm{Aggregate}_{{\bm{\theta}}}^{l}\left(\{h_{u}^{l-1},\forall u\in{\mathcal{N}}_{v}\}\right),$ (14)
where the neighborhood of $v\in V$ is denoted ${\mathcal{N}}_{v}=\{u\in V\mid(v,u)\in E\}$, ${\bm{h}}_{u}^{l-1}$ is the node feature of the neighborhood, and ${\bm{a}}_{v}^{l}$ is the aggregated node feature of the neighborhood. Then, the combined step at each $l$-th layer is defined as follows:
${\bm{h}}_{v}^{l}=\mathrm{Combine}^{l}_{{\bm{\theta}}}({\bm{h}}_{v}^{l-1},{\bm{a}}_{v}^{l}),$ (15)
where ${\bm{h}}_{v}^{l}\in{\mathbb{R}}^{H_{l}}$ denotes the node representation at the $l$-th layer. Here, the hyperparameters for the total number of layers $L$ and the intermediate vector dimension $N^{l}$ are determined empirically. Although numerous implementations of GNN architectures have been proposed to date, the most basic and widely used architecture is the GCN [Scarselli et al., 2008], which is given as follows:
${\bm{h}}_{v}^{l}=\sigma\left(W^{l}\sum_{u\in{\mathcal{N}}(v)}\frac{{\bm{h}}_{u}^{l-1}}{|{\mathcal{N}}(v)|}+B^{l}{\bm{h}}_{v}^{l-1}\right),$ (16)
where $W^{l}$ and $B^{l}$ are trainable parameters, $|{\mathcal{N}}(v)|$ serves as a normalization factor, and $\sigma:{\mathbb{R}}^{H_{l}}\to{\mathbb{R}}^{H_{l}}$ is some component-wise nonlinear activation function, e.g., the sigmoid or ReLU function.
Table 3: ApR for MaxCut on Gset
GRAPH (NODES, EDGES) GREEDY SDP RUN-CSP PI-GNN CRA
G1 ($800$, $19{,}176$) $0.942$ $0.970$ $0.979$ $0.978$ $\mathbf{1.000}$
G2 ($800$, $19{,}176$) $0.951$ $0.970$ $0.981$ $0.976$ $\mathbf{0.998}$
G3 ($800$, $19{,}176$) $0.945$ $0.972$ $0.982$ $0.972$ $\mathbf{1.000}$
G4 ($800$, $19{,}176$) $0.949$ $0.971$ $0.980$ $0.978$ $\mathbf{0.999}$
G5 ($800$, $19{,}176$) $0.949$ $0.970$ $0.980$ $0.978$ $\mathbf{1.000}$
G14 ($800$, $4{,}694$) $0.946$ $0.952$ $0.956$ $0.988$ $\mathbf{0.994}$
G15 ($800$, $4{,}661$) $0.939$ $0.958$ $0.952$ $0.980$ $\mathbf{0.992}$
G16 ($800$, $4{,}672$) $0.948$ $0.958$ $0.953$ $0.965$ $\mathbf{0.990}$
G17 ($800$, $4{,}667$) $0.946$- $0.956$ $0.967$ $\mathbf{0.990}$
G22 ($2{,}000$, $19{,}990$) $0.923$- $0.972$ $0.987$ $\mathbf{0.998}$
G23 ($2{,}000$, $19{,}990$) $0.927$- $0.973$ $0.968$ $\mathbf{0.997}$
G24 ($2{,}000$, $19{,}990$) $0.927$- $0.973$ $0.959$ $\mathbf{0.998}$
G25 ($2{,}000$, $19{,}990$) $0.929$- $0.974$ $0.974$ $\mathbf{0.998}$
G26 ($2{,}000$, $19{,}990$) $0.924$- $0.974$ $0.965$ $\mathbf{0.998}$
G35 ($2{,}000$, $11{,}778$) $0.942$- $0.953$ $0.968$ $\mathbf{0.990}$
G36 ($2{,}000$, $11{,}766$) $0.942$- $0.953$ $0.966$ $\mathbf{0.991}$
G37 ($2{,}000$, $11{,}785$) $0.946$- $0.950$ $0.966$ $\mathbf{0.997}$
G38 ($2{,}000$, $11{,}779$) $0.943$- $0.949$ $0.966$ $\mathbf{0.991}$
G43 ($1{,}000$, $9{,}990$) $0.928$ $0.968$ $0.976$ $0.966$ $\mathbf{0.995}$
G44 ($1{,}000$, $9{,}990$) $0.920$ $0.955$ $0.978$ $0.968$ $\mathbf{0.998}$
G45 ($1{,}000$, $9{,}990$) $0.930$ $0.950$ $0.979$ $0.961$ $\mathbf{0.998}$
G46 ($1{,}000$, $9{,}990$) $0.930$ $0.960$ $0.976$ $0.974$ $\mathbf{0.998}$
G47 ($1{,}000$, $9{,}990$) $0.931$ $0.956$ $0.976$ $0.972$ $\mathbf{0.997}$
G48 ($3{,}000$, $6{,}000$) $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{1.000}$ $0.912$ $\mathbf{1.000}$
G49 ($3{,}000$, $6{,}000$) $\mathbf{1.000}$ $\mathbf{1.000}$ $\mathbf{1.000}$ $0.986$ $\mathbf{1.000}$
G50 ($3{,}000$, $6{,}000$) $\mathbf{1.000}$ $\mathbf{1.000}$ $0.999$ $0.990$ $\mathbf{1.000}$
G51 ($1{,}000$, $5{,}909$) $0.949$ $0.960$ $0.954$ $0.964$ $\mathbf{0.991}$
G52 ($1{,}000$, $5{,}916$) $0.944$ $0.957$ $0.955$ $0.961$ $\mathbf{0.990}$
G53 ($1{,}000$, $5{,}914$) $0.945$ $0.956$ $0.950$ $0.966$ $\mathbf{0.992}$
G54 ($1{,}000$, $5{,}916$) $0.900$ $0.956$ $0.952$ $0.970$ $\mathbf{0.988}$
G55 ($5{,}000$, $12{,}498$) $0.892$- $0.978$ $0.983$ $\mathbf{0.991}$
G58 ($5{,}000$, $29{,}570$) $0.945$- $0.980$ $0.966$ $\mathbf{0.989}$
G60 ($7{,}000$, $17{,}148$) $0.889$- $0.980$ $0.945$ $\mathbf{0.991}$
G63 ($7{,}000$, $41{,}459$) $0.948$- $0.947$ $0.968$ $\mathbf{0.989}$
G70 ($10{,}000$, $99{,}9$ 999) $0.886$- $0.970$ $0.982$ $\mathbf{0.992}$
Appendix F Additional experiments
F.1 Numerical validation of practical issues presented in Section 3.1
In this sectioin, we will examine the issue (I) with continuous relaxations and the issue (II), the difficulties of optimization, as pointed out by previous studies [Wang and Li, 2023, Angelini and Ricci-Tersenghi, 2023], in the NP-hard problems of MIS and the MaxCut problem.
Therefere, we conducted numerical experiments using the PI-GNN solver for MIS and MaxCut problems on RRGs with higher degrees. To ensure the experimental impartiality, we adhered to the original settings of the PI-GNN solver [Schuetz et al., 2022b]. Refer to Section E for the detailed experimental settings. Fig. 7 (top) shows the solutions obtained by the PI-GNN solver as a function of the degree $d$ for the MIS and MaxCut problems with varying system sizes $N$. These results indicate that finding independent and cut sets becomes unfeasible as the RRG becomes denser. In addition, to clarify the reasons for these failures, we analyzed the dynamics of the cost function for MIS problems with $N=10{,}000$, with a specific focus on a graph with degrees $d=5$ and $d=20$, as depicted in Fig.  7 (bottom). For the $d=5$ case, the cost function goes over the plateau of $\hat{l}({\bm{\theta}};G,{\bm{\lambda}})=0$ with ${\bm{p}}_{\theta}(G)={\bm{0}}_{N}$, as investigated in the histogram, eventually yielding a solution comparable to those presented by Schuetz et al. [2022a]. Conversely, in the $d=20$ case, the cost function remains stagnant on the plateau of $\hat{l}({\bm{\theta}};G,{\bm{\lambda}})=0$ with ${\bm{p}}_{\theta}(G)={\bm{0}}_{N}$, thereby failing to find any independent nodes. Interpreting this phenomenon, we hypothesize that the representation capacity of the GNN is sufficiently large, leading us to consider the optimization of $\hat{L}_{\mathrm{MIS}}({\bm{\theta}};G,\lambda)$ and $\hat{L}_{\mathrm{MaxCut}}({\bm{\theta}};G)$ as a variational optimization problem relative to ${\bm{p}}_{\theta}$. In this case, ${\bm{p}}_{\theta}^{\ast}={\bm{0}}_{N}$ satisfies the first-order variational optimality conditions $\delta\hat{l}_{\mathrm{MIS}}/\delta{\bm{p}}_{\theta}|_{{\bm{p}}_{\theta}={\bm{p}}^{\ast}}=\delta\hat{l}_{\mathrm{MaxCut}}/\delta{\bm{p}}_{\theta}|_{{\bm{p}_{\theta}}={\bm{p}}^{\ast}}={\bm{0}}_{N}$, which implies a potential reason for absorption into the plateau. However, this does not reveal the conditions for the convergence to the fixed point ${\bm{p}}^{\ast}$ during the early learning stage or the condition to escape from the fixed point ${\bm{p}}^{\ast}$. Thus, an extensive theoretical evaluation through stability analysis remains an important topic for future work.
Figure 7: The top graph shows the independent set density for MIS problems (left) and the cut ratio for MaxCut problems (right) as a function of degree $d$ using the PI-GNN solver with varying system size $N$. Each data point represents the average result of five different graph instances, with the error bars indicating the standard deviation of those results. The bottom graph shows the cost as a function of the number of parameter updates $N_{\mathrm{EPOCH}}$, for $N=10000$ MIS problems on $5$-RRG and $20$-RRG. The histogram represents the relaxed vector distribution with varying numbers of parameter updates $N_{\mathrm{EPOCH}}$. Each point in the bottom-left plot is linked to the corresponding bottom-right histogram.
In summary, UL-based solver, minimizing ${\bm{\theta}}$ can be challenging and unstable. In particular, the PI-GNN solver, which is one of the UL-based solvers employing GNNs, fails to optimize ${\bm{\theta}}$ due to a local solution in complex CO problems on relatively dense graphs where the performance of greedy algorithms worsens. This issues can be potential bottleneck for more practical and relatively dense problems, making it challenging to employ the PI-GNN solver confidently.
F.2 Additional results of Gset
We conducted experiments across the additional GSET collection to further validate that including CRA enhances PI-GNN results beyond previously achievable in Table 3.
F.3 Ablation over initial scheduling value and scheduling rate
We conducted an ablation study focusing on the initial scheduling value $\gamma(0)$ and scheduling rate $\varepsilon$. This numerical experiment was conducted under the configuration described in Section 5 and E. Fig.  8 shows the IS density of $N=10000$ MIS problems on a $100$-RRG as a function of the initial scheduling value $\gamma(0)$ and the scheduling rate $\varepsilon$ using the CRA-PI-GNN with both GraphSage and GCV. As can be seen, smaller initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values typically yield better solutions. However, the convergence time increases progressively as the initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values become smaller. In addition, GraphSage consistently produces better solutions even with relatively larger initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values, which implies that the GNN architecture influences both the solution quality and the effective regions of the initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values for the annealing process.
Figure 8: (Top) IS density of $N=10{,}000$ MIS problems on $100$-RRG as a function of initial scheduling $\gamma(0)$ and scheduling rate $\varepsilon$ values obtained by the CRA-PI-GNN solver using GraphSage (Left) and GCV (Right). The color of the heat map represents the average IS over five different instances.
F.4 Ablation over curve rate
Next, we investigated the effect of varying the curvature $\alpha$ in Eq. 5. Numerical experiments were performed on MIS problems with $10{,}000$ nodes and the degrees of $5$ and $20$, as well as MaxCut problems with $10{,}000$ nodes and the degrees of $5$ and $35$. The GraphSAGE architecture was employed, with other parameters set as Section  5 adn E. As shown in Fig. 9, we observed that $\alpha=2$ consistently yielded the best scores across these problems.
Figure 9: (Left) Independent set density as a function of curveture rate $\alpha$ in Eq. (5). (Right) Cut ratio as a function of curveture rate $\alpha$ in Eq. (5). Error bars represent the standard deviations of the results